OpenStudy (anonymous):

I need help an investor invested 1,200 in two funds. One fund earned a 7% profit and one earned a 2% profit. The total profit is 39$ so how much was invested in each acount

OpenStudy (anonymous):

T + S = $1,200 .02T + .07S = $ 39 Can you solve it now?

OpenStudy (anonymous):

let me see I might be a minute

OpenStudy (anonymous):

.07+.02=.09, this is where I am stuck, finding the amount in the first acount

OpenStudy (anonymous):

2T + 7S = 3900 How about now?

OpenStudy (anonymous):

t+1950-7s/2

OpenStudy (anonymous):

How about now? 2 ( 1,200 - S ) + 7S = 3900

OpenStudy (anonymous):

what is the 3900 for?

OpenStudy (anonymous):

dollars

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

S = ? , T = ?

OpenStudy (anonymous):

t=2

OpenStudy (anonymous):

Do you know how to multiply? 2 ( 1,200 - S ) + 7S = 3900

OpenStudy (anonymous):

yes but I have been stuck on this one for 3 hours. I see how you got 3900

OpenStudy (anonymous):

I am missing part of the equation

OpenStudy (anonymous):

which part are you missing???

OpenStudy (anonymous):

I cannot figure out how to come up with the amount the fund that has 9%.. Can you show me what step I am missing?

OpenStudy (anonymous):

There's nothing 9% !!!

OpenStudy (anonymous):

how much was invested into the fund that had 7%. How much was invested into the fund that had 2% This is where I am lost. 1200 is split between 2 acounts

OpenStudy (anonymous):

I can't figure this part out

OpenStudy (anonymous):

Do you know how to multiply to obtain the amount 7%: 2 ( 1,200 - S ) + 7S = 3900

OpenStudy (anonymous):

yes I used a calculator

OpenStudy (anonymous):

Nothing to use calculator here!

OpenStudy (anonymous):

3900 is the amount of profit

OpenStudy (anonymous):

Are you capable of multiply by 2 * 1,200?

OpenStudy (anonymous):

2*1,200=2400

OpenStudy (anonymous):

Then, do it: 2 ( 1,200 - S ) + 7S = 3900

OpenStudy (anonymous):

(1,200-S) looses me

OpenStudy (anonymous):

Thank you very much for your patience and help. I need to leave this one alone for awhile.

OpenStudy (anonymous):

I will save this so I can review

OpenStudy (anonymous):

2 ( 1,200 - S ) + 7S = 3,900 2,400 - 2s + 7s = 3,900 -> s = ( 3,900 - 2,400) / 5 = 300 => T = ...?

OpenStudy (anonymous):

Good luck with your Algebra !!!